# 100,000 Crab giant pulses at 215 MHz detected with an SKA-Low prototype station

### v1

@@ Acknowledgments
EDA2 is hosted by the MWA under an agreement via the MWA External Instruments Policy. This scientific work makes use of the Murchison Radio-astronomy Observatory, operated by CSIRO. We acknowledge the Wajarri Yamatji people as the traditional owners of the Observatory site. This work was further supported by resources provided by the Pawsey Supercomputing Centre with funding from the Australian Government and the Government of Western Australia. The acquisition system was designed and purchased by INAF/Oxford University and the RX chain was design by INAF, as part of the SKA design and prototyping program. SD acknowledges the funding received from the ICRAR Summer Student program, which made parts of this work possible.
@@ Appendix B DM verification by maximising peak flux of GPs
It was noticed that high scattering may impact accuracy of timing-based DM measurements (mainly by shifting and changing shape of the average pulse profiles). Therefore, in order to verify the DM measurements obtained from timing we also measured DM using very bright (SNR20) individual GPs. We tested the following methods of finding the optimal DM by : (a) maximising the slope of the leading edge between and of the peak, (b) maximising the slope of the entire leading edge, (c) maximising the peak flux density (i.e. SNR) and (d) maximising the rise time (difference between the peak time and start of the leading edge). The first 2 methods (a,b) were significantly affected by the scattering leading to large errors and variations of the resulting DM during high-scattering period. Therefore, we rely on method (c), which is also a standard way used by single pulse search software packages finding pulse DM which maximises SNR/peak flux. Method (d) gives very similar results to method (c), but was also less reliable during high-scattering times. The comparison of DM measured with standard pulsar timing and by maximising peak flux density (method (c)) is shown in Figure 16. The errors in method (c) are much larger than from timing analysis during high scattering (between 10 and 28 February 2025) mainly because method (c) relies on single GPs (highly scattered and low-SNR at times), while timing analysis uses entire (1-hour) datasets. However, despite signs of systematic offset (especially between our measurements and Jodrell Bank data) the two methods are consistent, and indicate that the increase in DM, and its positive correlation with scattering time are real effects, and analysis artefacts caused by the scatter broadening. This comparison, as well as comparison with DM measured by Jodrell Bank gave us confidence that our DM measurements are to within a factor of 2 accurate.

### v2

@@ Acknowledgments
EDA2 is hosted by the MWA under an agreement via the MWA External Instruments Policy. This scientific work makes use of the Murchison Radio-astronomy Observatory, operated by CSIRO. We acknowledge the Wajarri Yamatji people as the traditional owners of the Observatory site. This work was further supported by resources provided by the Pawsey Supercomputing Centre with funding from the Australian Government and the Government of Western Australia. The acquisition system was designed and purchased by INAF/Oxford University and the RX chain was design by INAF, as part of the SKA design and prototyping program. SD acknowledges the funding received from the ICRAR Summer Student program, which made parts of this work possible. We thank the anonymous referee for their helpful comments that significantly improved the quality of the manuscript.
@@ Appendix B DM verification by maximising peak flux of GPs
It was noticed that high scattering may impact accuracy of timing-based DM measurements (mainly by shifting and changing shape of the average pulse profiles). Therefore, in order to verify the DM measurements obtained from timing we also measured DM using very bright (SNR20) individual GPs. We tested the following methods of finding the optimal DM by : (a) maximising the slope of the leading edge between and of the peak, (b) maximising the slope of the entire leading edge, (c) maximising the peak flux density (i.e. SNR) and (d) maximising the rise time (difference between the peak time and start of the leading edge). The first two methods (a,b) were significantly affected by the scatter-broadening which led to large errors and variations of the resulting DM during high-scattering period. Therefore, we rely on method (c), which is also the typically used by single pulse search software packages finding DM maximising SNR/peak flux. The method (d) gives very similar results to method (c), but was also less reliable during high-scattering times. The comparison of DM measured with standard pulsar timing and by maximising peak flux density (method (c)) is shown in Figure 17. The errors in method (c) are much larger than from timing analysis during high scattering (between 10 and 28 February 2025) mainly because method (c) relies on single GPs (highly scattered and low-SNR at times), while timing analysis uses entire (1-hour) datasets. However, despite signs of systematic offset (especially between our measurements and Jodrell Bank data) the two methods are consistent, and indicate that the increase in DM, and its positive correlation with scattering time are real effects (i.e. not artefacts of analysis caused by the scatter broadening). This comparison, as well as comparison with DM measured by Jodrell Bank gave us confidence that our DM measurements are to within a factor of 2 accurate.